myininaya (myininaya):

Find two nonzero values for x and y such that 15x-21y=3

OpenStudy (anonymous):

15x-21y=3 devide both sodes by 3 5x - 7y =1 add 7y both side 5x = 1+7y devide both sodes by 5 x = (1+7y)/5 let y= any nonzero number...say 2 for example then x = (1+ 7*2)/5 = 15/5 = 3 i.e x=3, y=2

myininaya (myininaya):

very good this wasn't the approach i was expecting lol i was expecting to see the euclidean alg. Nice Job!